OpenStudy (anonymous):

A football team can sell 10,000 tickets at $10 each. For each increase in price of $.25, the demand for tickets decreases by 200 tickets. How much should the team charge per ticket to maximize the revenue? (Only a calculus solution with appropriate derivative testing is acceptable) How much is the maximum revenue and how many tickets were sold to reach that maximum revenue?

jimthompson5910 (jim_thompson5910):

same idea as last time, just different numbers starting at $10 a ticket, we have 10,000 people that want them and are willing to pay for them

jimthompson5910 (jim_thompson5910):

as we increase the price, the amount of people that want them drops by 200 each time

jimthompson5910 (jim_thompson5910):

if you increase the price x times, then your final price will be y = 0.25x + 10 at the same time, 200x people will leave if you increase the price x times so you go from 10,000 people to 10000 - 200x people

jimthompson5910 (jim_thompson5910):

Revenue = (# of tickets sold)*(price per ticket) R(x) = (10000 - 200x)(0.25x + 10) R(x) = 10000(0.25x + 10) - 200x(0.25x + 10) R(x) = 2500x + 100000 - 50x^2 - 2000x R(x) = -50*x^2+500x+100000

jimthompson5910 (jim_thompson5910):

let me know when you want me to continue

OpenStudy (anonymous):

Since it is the same general idea I can sort of follow a little faster, it seems fairly similar my problem just seems to be creating the equations to begin with

jimthompson5910 (jim_thompson5910):

yeah it's all the same idea, just different numbers...that's all that's new

jimthompson5910 (jim_thompson5910):

I think they just wanted you to practice with this sort of thing...or just give you busy work lol

jimthompson5910 (jim_thompson5910):

so when you derive R(x) with respect to x, what do you get? R(x) = -50x^2+500x+100000 R ' (x) = ??

OpenStudy (anonymous):

-100x+500

jimthompson5910 (jim_thompson5910):

good, now set that equal to zero and solve for x

jimthompson5910 (jim_thompson5910):

to find your critical value

OpenStudy (anonymous):

okay, I got 5!

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

so if they increase the price exactly 5 times, they will max out the revenue

jimthompson5910 (jim_thompson5910):

the tickets will be y = 0.25x + 10 y = 0.25(5) + 10 y = 11.25 dollars each

jimthompson5910 (jim_thompson5910):

and there will be y = 10000 - 200x y = 10000 - 200(5) y = 9000 people that will buy the tickets

jimthompson5910 (jim_thompson5910):

this will make the max revenue to be R(x) = -50x^2+500x+100000 R(5) = -50(5)^2+500(5)+100000 R(5) = 101,250 dollars

OpenStudy (anonymous):

okay and I already was able to answer the last question! I'll post the last one and then we will be on our way!!